OpenStudy (anonymous):

Given f(x) show below, complete the equation for the inverse of F(x). If necessary, use the slash (/) for the division symbol. F(x)=8x f^-1(y)=____

OpenStudy (zzr0ck3r):

you have y = 8x solve for x x=y/8 so the inverse is f^(-1)(y) = y/8

OpenStudy (anonymous):

so x/8?

OpenStudy (zzr0ck3r):

they didn't switch the variables (and that is fine)

OpenStudy (zzr0ck3r):

so y/8, is the notation they are using.

OpenStudy (anonymous):

okay so its y/8 nit x/8 right

OpenStudy (zzr0ck3r):

some people like to switch the variables and solve for y, but its really not proper...

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

they told us our function is a function of y "f^-1(y)=____"

OpenStudy (anonymous):

ok got thanks im going to open another post and im going to try and do it can u check my work please

OpenStudy (debbieg):

It is typical to switch x & y, then solve for y, so both the original function and the inverse are of the form y= a function of x. But yeah, looks like here your teacher wants you to simply solve the original function, where you have y as a function of x, for x, so you get x and a function of y. A little bit odd, but it works. As long as you don't want to graph the function and the inverse together, lol.

OpenStudy (zzr0ck3r):

its actually odd to switch the variables, because then you are switching the domains. and if the function went from x to y, then the inverse should go from y to x

OpenStudy (zzr0ck3r):

but its common to switch them to make students understand the process but if f(x):A to B then f^(-1)(x) goes from B to A (but with this notation it is assumed that x is in A for the function, and x is in B for the inverse)